OpenStudy (anonymous):

John is going to rent a tent to go camping. The cost to rent a camping tent from Woodland Outfitters is a flat rate of $12.00 plus $2.00 for each night the tent is rented. Part a. Let x represent the number of nights John keeps the tent and write a cost function, c(x), for the cost to rent a tent from Woodland Outfitters. Part b. How much will John owe if he uses the tent for three nights?

OpenStudy (anonymous):

C(x) = 2x + 12 for part B, you'll need to calculate C(3)... C(3) = 2(3) + 12 = ??? can you take it from here?

OpenStudy (anonymous):

i get the first part but im still confused on the second part. Thanks thought! :)

OpenStudy (anonymous):

2*3 + 12 = 6 + 12 = ???

OpenStudy (anonymous):

yw...:)